# Tiling spaces are covering spaces over irrational tori

### v3

@@ Abstract
We study tiling spaces in the diffeological context. We prove some basic diffeological properties for tiling spaces and analyze two different fiber bundle structures of one-dimensional tiling spaces over the irrational torus. We use the diffeological classification of irrational tori which captures their arithmetical escence in order to inherit the diffeological equivalence in the context of tiling spaces.
@@ 1 Introduction
Our motivation comes from the naive question: what is the aperiodic analog of the orbifolds that characterize the symmetries of periodic tilings of the plane? Conway’s Magic Theorem [9] classifies the symmetries of periodic tilings of the plane through the orbifolds obtained as quotients $\mathcal{O}=\mathbb{R}^{2}/G$, where $G$ is the symmetry group of the tiling. In the crystallographic case, $G$ has a subgroup of translations isomorphic to $\mathbb{Z}^{2}$. Thus the aforementioned orbifolds are quotients of the torus $\mathbb{T}^{2}$ by a discrete subgroup of $G$. In the aperiodic one-dimensional case, however, one will find that a plausible analog of the translations subgroups $\mathbb{Z}$ could be (a subset of) $\mathbb{Z}+\alpha\mathbb{Z}$, with $\alpha$ an irrational number. So one should look at $\mathbb{T}_{\alpha}:=\mathbb{R}/(\mathbb{Z}+\alpha\mathbb{Z})$, which is quite a strange object, its natural topology is coarse. It is nevertheless non-trivial when endowed with a structure of diffeological space, it is called the irrational torus of slope $\alpha$. It was first studied by Iglesias-Zemmour and Donato in [20], see also [21, 18]. On the other hand, another possible analog for $\mathcal{O}$ could be the tiling space $\Omega$, also known as the continuous hull of a tiling. For instance, the tiling space of a periodic tiling is a torus obtained by identifying translations. If further symmetries are to be considered then one will end up with Conway’s orbifolds.
@@ 2.1.2 Tiling spaces
$B_{\frac{1}{\varepsilon}}(0)\cap T+x=B_{\frac{1}{\varepsilon}}(0)\cap T^{\prime}+x^{\prime}.$
@@
A substitution rule $\omega$ is called primitive if every proto-tile appears in some iteration of the substitution rule applied to any other proto-tile. Primitive substitutions generates unique $LI$ classes, thus every tiling generate the same hull and so we may speak about a tiling space $\Omega_{\omega}$ associated to $\omega$.
@@
Figure 1: The branched 1-dimensional manifolds $\Omega_{0}$, $\Omega_{1}$ and $\Omega_{2}$ associated to the subtitution rule $a\mapsto b$, $b\mapsto ab$. Note that this image is only for illustrative purposes but it is not precisely what is needed – we are omitting here a technicality called forcing the border (see Theorem 4.3 in [2]).
@@
Tiling spaces are fiber bundles over torus with totally disconnected fiber.
@@ 2.1.3 Equivalence of tiling spaces
The first kind of equivalence between tiling spaces that we will encounter in sintony with diffeological equivalences is the following:
@@ 2.1.4 Geometrical and combinatorial aspects of tiling spaces
$\alpha=\frac{c+\beta d}{a+\beta b}=M\beta\quad\text{with}\quad ad-bc=\pm 1.$ (4)
@@ 2.2.1 Diffeological Spaces
For a smooth surjection $\pi:X\to B$ between diffeological spaces $X,B$ it is possible to define the structure groupoid $K_{\pi}$ of the surjection $\pi$ as the groupoid whose objects are $\mathrm{Obj}(K_{\pi})=B$ and its arrows are $\mathrm{Mor}(K_{\pi})=\bigcup_{b,b^{\prime}\in B}C^{\infty}(X_{b},X_{b^{\prime}})$, where $X_{b}$ denotes the fiber $\pi^{-1}(b)$. See [19, Postulate 8.2].
@@
The existence of the $D$-topology allows us to talk about local diffeological properties, defined in $D$-neighborhoods. A map $f:X\to X^{\prime}$ between diffeological spaces is said to be locally smooth in $x\in X$ if there is a $D$-open subset $A$ containing $x$ such that the map $f|_{A}$ is smooth. That is, if for every plot $P$ of $X$ the parametrization $f\circ P|_{P^{-1}(A)}$ is a plot of $X^{\prime}$. The map $f:X\to X^{\prime}$ is smooth if and only if it is locally smooth in every point of $X$. See [19, Postulate 2.3].
@@ 2.2.2 Diffeological Fiber Bundles
[19, Postulate 8.9] Let $X$ and $B$ diffeological spaces. A smooth map $\pi:X\to B$ is a fibration if and only if there exists a diffeological space $F$ such that the pullback $P^{*}(X)$ of $\pi$ along any plot $P:U\to B$ of $B$ is locally trivial with fiber $F$, where
@@
[19, Postulate 8.11] Consider the projection $\pi:X\to X/G$. If $\mathcal{A}_{X}^{G}$ is an induction then $\pi$ is a diffeological fibration with fiber $G$ and the action of $G$ on $X$ is called principal. Moreover, any equivalent fiber bundle $\pi^{\prime}:X\to Q$ to $\pi$ is a $G$-principal fiber bundle or $G$-principal fibration.
@@ 2.2.3 The irrational torus
is a diffeological space and it is diffeomorphic to $\mathbb{T}^{2}/\mathcal{S}_{\alpha}$ (see [18, Proposition 3]). The density of $\mathbb{Z}+\alpha\mathbb{Z}$ in $\mathbb{R}$ implies that $\pi_{\alpha}$ is locally trivial along the plots of $\mathbb{T}_{\alpha}$ (see Lemma 2) but it is not locally trivial.
@@
For a more complete treatment we refer the reader to [18, Propositions 4-7].
@@
Another property worth mentioning is that the exact sequence of homotopy for fiber bundles also holds in the diffeological context, (see [19, Postulate 8.21]). In our case, it gives us the following exact sequence:
@@ 3.1 The Diffeological Space $\Omega$
and $\mathcal{D}_{1}=\langle F\rangle$. We have that the diffeology $\mathcal{D}$ from Definition 12 contains $\mathcal{D}_{1}$. But $\mathcal{D}$ is minimal by definition of generating family. Thus $\mathcal{D}=\mathcal{D}_{1}$.
By definition of generating family [19, Postulate 1.68], the plots of $\mathcal{D}=\langle F\rangle$ are specified as the parametrizations $P:U\to\Omega$ such that for every $u\in U$ there exists an open neighborhood $V\subset U$ of $u$ such that either $P|_{V}$ is a constant parametrization or there exists an element $\hat{T}:\mathbb{R}^{d}\to\Omega$ of $\mathcal{F}$ and a smooth map $Q:V\to\mathbb{R}^{d}$ such that $P|_{V}=\hat{T}\circ Q$. As shown in the following diagram.
@@ 4 Diffeological and sturmian equivalences for Tiling spaces
Note that the equivalence classes $[x]$ of $\mathbb{T}^{2}/\mathcal{S}_{\alpha}$ correspond to orbits under the action by translation on $\Omega_{\alpha}$, so they are path connected and diffeomorphic to $\mathbb{R}$. We define the irrational tori with two origins $\mathbb{T}_{\alpha}^{\bullet}$ as the quotient of the punctured torus $\mathbb{T}^{2}-\{e\}$ by the equivalence relation $x\sim y$ if and only if $[x]=[y]$ and both $x$ and $y$ are in the same path connected component of $[x]-\{e\}$, for $e\in\mathcal{S}_{\alpha}$ the identity element of the group $\mathbb{T}^{2}$. The class $\mathcal{S}_{\alpha}$ is divided in two rays, which are diffeomorphic to the real line.
Once this construction is made, it is straightforward to check that the quotient map $\pi:\mathbb{T}^{2}-\{e\}\to\mathbb{T}_{\alpha}^{\bullet}$ has fibrating structure groupoid $K_{\pi}$. So it is a diffeological fibration with fiber $\mathbb{R}$. The inclusion $i:\mathbb{T}^{2}-\{e\}\to\mathbb{T}^{2}$ induces the following conmutative diagram:
[…]
note that $I_{\bullet}$ is surjective. Thus, we define the $\mathbb{Z}$-principal fiber bundle $X_{\alpha}\to\mathbb{T}_{\alpha}^{\bullet}$
Now we can extend through $\Psi$ in order to obtain the following conmutative diagram of smooth maps
[…]
@@
The proof relies on the exact sequence of homotopy [19, Postulate 8.21] of the diffeological fibration $\pi:\mathbb{T}^{2}-\{e\}\to\mathbb{T}_{\alpha}^{\bullet}$,

### v4

@@ Abstract
We study tiling spaces in the diffeological context. We prove some basic diffeological properties for tiling spaces and analyze two different fiber bundle structures of tiling spaces over irrational tori. We use the diffeological classification of irrational tori which captures their arithmetical escence in order to inherit the diffeological equivalence in the context of one-dimensional tiling spaces.
@@ 1 Introduction
Our motivation comes from the naive question: what is the aperiodic analog of the orbifolds that characterize the symmetries of periodic tilings of the plane? Conway’s Magic Theorem [9] classifies the symmetries of periodic tilings of the plane through the orbifolds obtained as quotients $\mathcal{O}=\mathbb{R}^{2}/G$, where $G$ is the symmetry group of the tiling. In the crystallographic case, $G$ has a subgroup of translations isomorphic to $\mathbb{Z}^{2}$. Thus the aforementioned orbifolds are quotients of the torus $\mathbb{T}^{2}$ by a discrete subgroup of $G$. In the aperiodic one-dimensional case, however, one will find that a plausible analog of the translations subgroups $\mathbb{Z}$ could be (a subset of) $\mathbb{Z}+\alpha\mathbb{Z}$, with $\alpha$ an irrational number. So one should look at $\mathbb{T}_{\alpha}:=\mathbb{R}/(\mathbb{Z}+\alpha\mathbb{Z})$, which is quite a strange object, its natural topology is coarse. It is nevertheless non-trivial when endowed with a structure of diffeological space, it is called the irrational torus of slope $\alpha$. It was first studied by Iglesias-Zemmour and Donato in [21] (see also [22, 19]). On the other hand, another possible analog for $\mathcal{O}$ could be the tiling space $\Omega$, also known as the continuous hull of a tiling. For instance, the tiling space of a periodic tiling is a torus obtained by identifying translations. If further symmetries are to be considered then one will end up with Conway’s orbifolds.
@@ 2.1.2 Tiling spaces
$B_{\frac{1}{\varepsilon}}(0)\cap(T+x)=B_{\frac{1}{\varepsilon}}(0)\cap(T^{\prime}+x^{\prime}).$
@@
A substitution rule $\omega$ is called primitive if there is an integer number $N$ such that every proto-tile appears in the $N$-th iteration of the substitution rule applied to every other proto-tile. Primitive substitutions generate unique $LI$ classes, thus every tiling generates the same hull and so we may speak about a tiling space $\Omega_{\omega}$ associated to $\omega$.
@@
Remark 1.
We omited here the technicality of whether the substitution rule forces the border — which essentially means that, by applying the substitution, the patches would “grow in every direction”, not as, for example, the chair tiling substitution which would typically grow over the first quadrant of the plane. When a substitution does not force the border, there is an additional step in the proof of the previous theorem called collaring, which results in a modified complex $\tilde{\Omega}_{0}$. See [36] for details.
Figure 1: The branched 1-dimensional manifolds $\Omega_{0}$, $\Omega_{1}$ and $\Omega_{2}$ associated to the subtitution rule $a\mapsto b$, $b\mapsto ab$. Note that this image is only for illustrative purposes but it is not precisely what is needed – we are omitting here a technicality called forcing the border, see Remark 1.
@@
Tiling spaces are fiber bundles over tori with totally disconnected fiber.
@@ 2.1.3 Equivalence of tiling spaces
The first kind of equivalence between tiling spaces that we will encounter in concordance with diffeological equivalences is the following:
@@ 2.1.4 Geometrical and combinatorial aspects of tiling spaces
$\alpha=\frac{c+\beta d}{a+\beta b}=M\beta.$ (4)
@@ 2.2.1 Diffeological Spaces
For a smooth surjection $\pi:X\to B$ between diffeological spaces $X,B$ it is possible to define the structure groupoid $K_{\pi}$ of the surjection $\pi$ as the groupoid whose objects are $\mathrm{Obj}(K_{\pi})=B$ and its arrows are $\mathrm{Mor}(K_{\pi})=\bigcup_{b,b^{\prime}\in B}C^{\infty}(X_{b},X_{b^{\prime}})$, where $X_{b}$ denotes the fiber $\pi^{-1}(b)$. See [20, Postulate 8.2].
@@
The existence of the $D$-topology allows us to talk about local diffeological properties, defined in $D$-neighborhoods. A map $f:X\to X^{\prime}$ between diffeological spaces is said to be locally smooth in $x\in X$ if there is a $D$-open subset $A$ containing $x$ such that the map $f|_{A}$ is smooth. That is, if for every plot $P$ of $X$ the parametrization $f\circ P|_{P^{-1}(A)}$ is a plot of $X^{\prime}$. The map $f:X\to X^{\prime}$ is smooth if and only if it is locally smooth in every point of $X$. See [20, Postulate 2.3].
@@ 2.2.2 Diffeological Fiber Bundles
[20, Postulate 8.9] Let $X$ and $B$ diffeological spaces. A smooth map $\pi:X\to B$ is a fibration if and only if there exists a diffeological space $F$ such that the pullback $P^{*}(X)$ of $\pi$ along any plot $P:U\to B$ of $B$ is locally trivial with fiber $F$, where
@@
[20, Postulate 8.11] Consider the projection $\pi:X\to X/G$. If $\mathcal{A}_{X}^{G}$ is an induction then $\pi$ is a diffeological fibration with fiber $G$ and the action of $G$ on $X$ is called principal. Moreover, any equivalent fiber bundle $\pi^{\prime}:X\to Q$ to $\pi$ is a $G$-principal fiber bundle or $G$-principal fibration.
@@ 2.2.3 The irrational torus
is a diffeological space and it is diffeomorphic to $\mathbb{T}^{2}/\mathcal{S}_{\alpha}$ (see [19, Proposition 3]). The density of $\mathbb{Z}+\alpha\mathbb{Z}$ in $\mathbb{R}$ implies that $\pi_{\alpha}$ is locally trivial along the plots of $\mathbb{T}_{\alpha}$ (see Lemma 2) but it is not locally trivial.
@@
For a more complete treatment we refer the reader to [19, Propositions 4-7].
@@
Another property worth mentioning is that the exact sequence of homotopy for fiber bundles also holds in the diffeological context, (see [20, Postulate 8.21]). In our case, it gives us the following exact sequence:
@@ 3.1 The Diffeological Space $\Omega$
and $\mathcal{D}_{1}=\langle F\rangle$. We have that the diffeology $\mathcal{D}$ from Definition 12 contains $\mathcal{D}_{1}$. But $\mathcal{D}$ is minimal by definition of generating family [20, Postulate 1.68]. Thus $\mathcal{D}=\mathcal{D}_{1}$.
By definition of generating family, the plots of $\mathcal{D}=\langle F\rangle$ are specified as the parametrizations $P:U\to\Omega$ such that for every $u\in U$ there exists an open neighborhood $V\subset U$ of $u$ such that either $P|_{V}$ is a constant parametrization or there exists an element $\hat{T}:\mathbb{R}^{d}\to\Omega$ of $\mathcal{F}$ and a smooth map $Q:V\to\mathbb{R}^{d}$ such that $P|_{V}=\hat{T}\circ Q$. As shown in the following diagram.
@@ 4 Diffeological and sturmian equivalences for Tiling spaces
Note that the equivalence classes $[x]$ of $\mathbb{T}^{2}/\mathcal{S}_{\alpha}$ correspond to orbits under the action by translation on $\Omega_{\alpha}$, so they are path connected and diffeomorphic to $\mathbb{R}$. We define the irrational tori with two origins $\mathbb{T}_{\alpha}^{\bullet}$ as the quotient of the punctured torus $\mathbb{T}^{2}-\{e\}$ by the equivalence relation $x\sim y$ if and only if $[x]=[y]$ and both $x$ and $y$ are in the same path connected component of $[x]-\{e\}$, for $e\in\mathcal{S}_{\alpha}$ the identity element of the group $\mathbb{T}^{2}$. The class $\mathcal{S}_{\alpha}$ is divided in two rays, which are diffeomorphic to the real line. We will show next that this quotient map is actually a diffeological fibration.
The inclusion $i:\mathbb{T}^{2}-\{e\}\to\mathbb{T}^{2}$ induces the following conmutative diagram of smooth maps:
${\mathbb{T}^{2}-\{e\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$$\scriptstyle{\pi}$${\mathbb{T}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{\alpha}}$${\mathbb{T}_{\alpha}^{\bullet}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i_{\bullet}}$${\mathbb{T}_{\alpha},}$ (19)
where $i_{\bullet}$ is surjective. Let $P:V\to\mathbb{T}_{\alpha}^{\bullet}$ be a plot, so $i_{\bullet}\circ P=:Q$ is a plot of $\mathbb{T}_{\alpha}$. As $\mathbb{T}^{2}$ is a $\mathbb{R}$-principal bundle over $\mathbb{T}_{\alpha}$, by Lemma 2 we know that, for every $v_{0}\in V$, there is an open subset $U\subset V$ such that $Q|_{U}^{*}(\mathbb{T}_{\alpha})$ is a trivial bundle. Which means that there is a diffeomorphism $\eta$
$\eta:U\times\mathbb{R}\to Q|_{U}^{*}(\mathbb{T}^{2}).$
Also note that the inclusion $i:\mathbb{T}^{2}-\{e\}\to\mathbb{T}^{2}$ induces an embedding between the pullbacks
$P|_{U}^{*}(\mathbb{T}^{2}-\{e\})\hookrightarrow Q|_{U}^{*}(\mathbb{T}^{2}).$
Next we consider the map
$\mathrm{id}\times\exp:$ $U\times\mathbb{R}\to U\times\mathbb{R}$
$(v,t)\mapsto(v,\exp(t))$
and let $\bar{\eta}=\eta\circ(\mathrm{id}\times\exp):U\times\mathbb{R}\to Q|_{U}^{*}(\mathbb{T}^{2})$. Note that the image $\bar{\eta}(U\times\mathbb{R})$ of $\bar{\eta}$ does not contain the neutral element $e$ and thus
$\bar{\eta}(U\times\mathbb{R})\subset P|_{U}^{*}(\mathbb{T}^{2}-\{e\})\subset Q|_{U}^{*}(\mathbb{T}^{2}).$
Now let $U\times J=U\times[a,b]\subset U\times\mathbb{R}$. We have that $\bar{\eta}(U\times J)\cap\mathrm{pr}_{1}^{-1}(v_{0})$, where $\mathrm{pr}_{1}$ denotes the projection $\mathrm{pr}_{1}:P|_{U}^{*}(\mathbb{T}^{2})\subset U\times(\mathbb{T}^{2}-\{e\})\to U$, is an interval, say $[a_{v_{0}},b_{v_{0}}]$. This means that for every $v\in U$ and every interval $[a,b]$ there is an interval $[a_{v},b_{v}]$ such that
$\bar{\eta}(U\times[a,b])\cap\mathrm{pr}_{1}^{-1}(v)=[a_{v},b_{v}].$
So we can exhibit sections $s_{a}$, $s_{b}:U\to P|_{U}^{*}(\mathbb{T}^{2}-\{e\})$ given by
$s_{a}(u)$ $=\bar{\eta}(u,a),$
$s_{b}(u)$ $=\bar{\eta}(u,b)$
which allow us to define linear diffeomorphisms $\gamma_{u}:\mathrm{pr}_{1}^{-1}(u)\to\mathbb{R}$, such that $\gamma_{u}(s_{a}(u))=0$ and $\gamma_{u}(s_{b}(u))=1$. Then there is a diffeomorphism
$\mathbf{r}:P|_{U}^{*}(\mathbb{T}^{2}-\{e\})$ $\to U\times\mathbb{R}$
$x$ $\mapsto(u,\gamma_{u}(x)),$
where $u=\mathrm{pr}_{1}(x)$. Which means that $\pi:\mathbb{T}^{2}-\{e\}\to\mathbb{T}_{\alpha}^{\bullet}$ is locally trivial along the plots of $\mathbb{T}_{\alpha}^{\bullet}$ and, by Lemma 2, it is a diffeological fibration.
Finally, using 18 and 19, we define the $\mathbb{Z}$-principal fiber bundle $X_{\alpha}\to\mathbb{T}_{\alpha}^{\bullet}$ which can be extended through $\Psi$ in order to obtain the following conmutative diagram of smooth maps.
${X_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi}$$\scriptstyle{p}$${\Omega_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{\alpha}^{\bullet}}$${\mathbb{T}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{\alpha}}$${\mathbb{T}_{\alpha}^{\bullet}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i_{\bullet}}$${\mathbb{T}_{\alpha}}$ (20)
@@
The proof relies on the exact sequence of homotopy [20, Postulate 8.21] of the diffeological fibration $\pi:\mathbb{T}^{2}-\{e\}\to\mathbb{T}_{\alpha}^{\bullet}$,